OpenStudy (anonymous):

How do you find iterates of a function? For example, two problems I have are: 1) Find the first three iterates of the function f(Z) = Z^2 + 2 + i , when Z base 0 = 2 + 2i. 2) Find the third iterate of the function f(t) = 9t - 3 when t base 0 = 1/3 + i/9.

OpenStudy (kc_kennylau):

\[Z_1=Z_0^2+2+i\]

OpenStudy (kc_kennylau):

\[Z_1=Z_0^2+2+i=(2+2i)^2+2+i\]

OpenStudy (loser66):

z1 = f(z0) z2 = f(z1) z3=f(z2) that's it. just compute one by one

OpenStudy (anonymous):

i would use a calculator

OpenStudy (anonymous):

\[z_1=(2+2i)^2+2+i=2+9i\] http://www.wolframalpha.com/input/?i=%282%2B2i%29^2%2B2%2Bi

OpenStudy (anonymous):

\(z_2=(2+9i)^2+2+i=-75+37i\) http://www.wolframalpha.com/input/?i=%282%2B9i%29^2%2B2%2Bi

OpenStudy (anonymous):

should be clear how to do the next one right?

OpenStudy (anonymous):

oooohhhhh, i see what you mean! :D Okay, yes, that makes sooo much sense. Thank you, thank you, thank you!! ♥

OpenStudy (anonymous):

yw \(\heartsuit\)